determine whether the function f(x)=x+1 is odd,even,or neither
Answer by HyperBrain(694)

f(x)=x+1
f(-x)=-x+1
Since f(x) and f(-x) are not equal then the function is not even.
***it is a pitfall to say that if a function is not even, then it is odd and vice-versa. We shall see that this statement is not applicable here.
-f(x)=-x-1
---
Since f(-x) and -f(x) are not equal then, the function is not odd.
)_____
Therefore, the function is neither odd nor even